(3a-5)(9a2+25)(3a+5)

1
Expert's answer
2022-03-11T03:07:21-0500
Multiply 3a-5 by 3a+5

Apply this law: (a+b)(a-b)=( a2-b2)


(3a+5)(3a-5)=(9a2-25)

Then multiply (9a2-25) by (9a2+25) applying the above law.


(9a2-25)(9a2+25)=(81a4-625)